OpenStudy (anonymous):

use logarithmic differentiation to find the derivative of the function. y=(x^3+2)^2(x^5+4)^4

terenzreignz (terenzreignz):

\[y=\left( x^{3}+2 \right)^{2}\left( x^{5}+4 \right)^{4}\] Is that right?

terenzreignz (terenzreignz):

While you don't technically have to use logarithmic differentiation here, evaluating the powers of these binomials and then multiplying said powers and only then differentiating normally... might be a pain.. so let's get started :) \[\ln y=\ln \left[ \left( x^{3}+2 \right)^{2}\left( x^{5}+4 \right)^{4} \right]\]

OpenStudy (anonymous):

I agree with your comment. Y use log?

OpenStudy (anonymous):

I was thinking of using implicit as well

terenzreignz (terenzreignz):

Following the properties of the logarithm: \[\ln y = \ln \left( x^{3}+2 \right)^{2}+\ln \left( x^{5}+4 \right)^{4}\] And yet more properties of logarithms... \[\ln y = 2\ln \left( x^{3}+2 \right)+4\ln \left( x^{5}+4 \right)\]

OpenStudy (anonymous):

So far so good. I get it yay

terenzreignz (terenzreignz):

Use log because... it has this awesome ability to turn exponents into factors :D Also to turn products into sums... something like that

terenzreignz (terenzreignz):

Can you work it out from there?

OpenStudy (anonymous):

Hmm not sure about what u just said. But sound awesome. Lol

OpenStudy (anonymous):

Please help some more

terenzreignz (terenzreignz):

Well look at it: the 2 and the 4 which were once exponents, now they are just constants multiplied to logs... also, those two binomials were multiplied to each other, but now they are the sum of logs :) If nothing else, use logs because the instructions say so :P

terenzreignz (terenzreignz):

Maybe later, I'll show you a more practical use of logs, but for now... See that last expression there? Differentiate it implicitly....

OpenStudy (anonymous):

What last expression thw one u sent earlier

OpenStudy (anonymous):

If u can provide answer. I will get my own and match it to urs. Btw thanks so much U r awesome

terenzreignz (terenzreignz):

So you should get \[\frac{1}{y}\frac{dy}{dx}=\left( \frac{2}{x^{3}+2} \right)\left( 3x^{2} \right)+\left( \frac{4}{x^{5}+4} \right)\left( 5x^{4} \right)\]

terenzreignz (terenzreignz):

Now it's easier to see the derivative... don't you agree?

terenzreignz (terenzreignz):

Anyway, to get the derivative, just take that expression there on the right, and multiply it to y, which is given as a function of x, as in my first comment. And you're done! Also, there are such functions, like \[y = x^{x}\] for which I don't know how to get the derivative except via logarithmic differentiation.

OpenStudy (anonymous):

Yes i got this as well. Yay but i c the reason behind log derivatives thanks again

terenzreignz (terenzreignz):

No sweat